OpenStudy (amenah8):

find dy/dx of y=e^(cos(x^2))

OpenStudy (cathyangs):

How far have you worked with this? My first step would be to say that dy/dx of y=e^ N is E^N dN/dx , where dN/dx is ofc the derivative of the power with respect to x. Following?

OpenStudy (cathyangs):

I think you're gone, but here's a video to help you with chain rule ( https://www.khanacademy.org/math/differential-calculus/taking-derivatives/chain-rule/v/chain-rule-definition-and-example) Basically, I would then find dN/dx, where N = cos (x^2), which takes chain rule.

OpenStudy (irishboy123):

\(\Large y=e^{cos(x^2)}\) \(\Large \ln (y)=cos(x^2)\) \(\Large \frac{1}{y} y'= -\sin(x^2).2x \) etc etc